OpenStudy (anonymous):

Using complete sentences, explain which method you would use to solve the following system of equations and why. In your answer, include the solution to one of the variables and how you found it using the method you chose. 2x + y + z = –7 x – 3y + 4z = –14 x – 2y – 3z = –11

OpenStudy (anonymous):

I DONT KNOW HOW TO DO THIS

OpenStudy (anonymous):

I would use the method of elimination subtract last equation from the second to eliminate x that will give you -y + 7z = -3.......(1) now multiply second equation by 2 and subtract it from first to give another equation in y and z so we get 2x-6y+8z = -28 subtract this from first equation to get 7y -7z = 21 y-z = 3 now add this to equation (1) to give 6z = 0 z = 0 so y = 3 and find x by plugging in z=0 and y=3 into the first of the origional equations

OpenStudy (anonymous):

this method is less complicated than the substitution method another method would be to use matrices but maybe you haven't come to study that yet.